\begin{thebibliography}{99}                                                                                               %


\bibitem {pelaezrev}J.~R.~Pelaez,
%``From controversy to precision on the sigma meson: a review on the status of the non-ordinary $f_0(500)$ resonance,''
arXiv:1510.00653 [hep-ph].
%%CITATION = ARXIV:1510.00653;%%


\bibitem {pdg}K. A. Olive et al. (Particle Data Group), Chin. Phys.
\textbf{C38}, 090001 (2014).

\bibitem {caprini}
%\cite{Caprini:2005zr}
%\bibitem{Caprini:2005zr}
I.~Caprini, G.~Colangelo and H.~Leutwyler,
%``Mass and width of the lowest resonance in QCD,''
Phys.\ Rev.\ Lett.\ \textbf{96} (2006) 132001.
%[hep-ph/0512364].
%%CITATION = HEP-PH/0512364;%%


\bibitem {kaminski}
%\cite{GarciaMartin:2011jx}
%\bibitem{GarciaMartin:2011jx}
R.~Garcia-Martin, R.~Kaminski, J.~R.~Pelaez and J.~Ruiz de Elvira,
%``Precise determination of the f0(600) and f0(980) pole parameters from a dispersive data analysis,''
Phys.\ Rev.\ Lett.\ \textbf{107} (2011) 072001;
%[arXiv:1107.1635 [hep-ph]];
%\cite{Kaminski:2006qe}
%\bibitem{Kaminski:2006qe}
R.~Kaminski, J.~R.~Pelaez and F.~J.~Yndurain,
%``The Pion-pion scattering amplitude. III. Improving the analysis with forward dispersion relations and Roy equations,''
Phys.\ Rev.\ D \textbf{77} (2008) 054015.
%[arXiv:0710.1150 [hep-ph]].


\bibitem {our}
%\cite{Broniowski:2015oha}
%\bibitem{Broniowski:2015oha}
W.~Broniowski, F.~Giacosa and V.~Begun,
%``Cancellation of the $\sigma$ meson in thermal models,''
Phys.\ Rev.\ C \textbf{92} (2015) no.3, 034905.
%doi:10.1103/PhysRevC.92.034905.
%[arXiv:1506.01260 [nucl-th]].
%%CITATION = doi:10.1103/PhysRevC.92.034905;%%


\bibitem {wfbook}
%\cite{Florkowski:2010zz}
%\bibitem{Florkowski:2010zz}
W.~Florkowski, \textquotedblleft Phenomenology of Ultra-Relativistic Heavy-Ion
Collisions,\textquotedblright\ Singapore, Singapore: World Scientific (2010)
416 p.

\bibitem {giorgio}
%\cite{Torrieri:2004zz}
G.~Torrieri et al,
%``SHARE: Statistical hadronization with resonances,''
Comput.\ Phys.\ Commun.\ \textbf{167} (2005) 229;
%[nucl-th/0404083];
%%CITATION = NUCL-TH/0404083;%%
%\cite{Alba:2014eba}
%\bibitem{Alba:2014eba}
P.~Alba et al,
%, S. Steinke, W. Broniowski, W. Florkowski, J. Letessier and J. Rafelski,
%W. Alberico, R. Bellwied, M. Bluhm, V. Mantovani Sarti, M. Nahrgang and C. Ratti``Freeze-out conditions from net-proton and net-charge fluctuations at RHIC,''
Phys.\ Lett.\ B \textbf{738} (2014) 305;
%[arXiv:1403.4903 [hep-ph]];
%%CITATION = ARXIV:1403.4903;%%
%\cite{Andronic:2005yp}
%\bibitem{Andronic:2005yp}
A.~Andronic, P.~Braun-Munzinger and J.~Stachel,
%``Hadron production in central nucleus-nucleus collisions at chemical freeze-out,''
Nucl.\ Phys.\ A \textbf{772} (2006) 167.
%[nucl-th/0511071].
%%CITATION = NUCL-TH/0511071;%%


\bibitem {andronic}
%\cite{Andronic:2008gu}
%\bibitem{Andronic:2008gu}
A.~Andronic, P.~Braun-Munzinger and J.~Stachel,
%``Thermal hadron production in relativistic nuclear collisions: The Hadron mass spectrum, the horn, and the QCD phase transition,''
Phys.\ Lett.\ B \textbf{673} (2009) 142 [Phys.\ Lett.\ B \textbf{678} (2009)
516].
%[arXiv:0812.1186 [nucl-th]].
%%CITATION = ARXIV:0812.1186;%%


\bibitem {dashen}
%\cite{Dashen:1969ep}
%\bibitem{Dashen:1969ep}
R.~Dashen, S.~K.~Ma and H.~J.~Bernstein,
%``S Matrix formulation of statistical mechanics,''
Phys.\ Rev.\ \textbf{187} (1969) 345;
%%CITATION = PHRVA,187,345;%%
%\cite{Dashen:1974jw}
%\bibitem{Dashen:1974jw}
R.~F.~Dashen and R.~Rajaraman,
%``Narrow Resonances in Statistical Mechanics,''
Phys.\ Rev.\ D \textbf{10} (1974) 694.
%%CITATION = PHRVA,D10,694;%%


\bibitem {weinhold}
%\cite{Weinhold:1996ts}
%\bibitem{Weinhold:1996ts}
W.~Weinhold, B.~L.~Friman and W.~Noerenberg,
%``Thermodynamics of an interacting pi N system,''
Acta Phys.\ Polon.\ B \textbf{27} (1996) 3249;
%%CITATION = APPOA,B27,3249;%%
%\cite{Weinhold:1997ig}
%\bibitem{Weinhold:1997ig}
W.~Weinhold, B.~Friman and W.~Norenberg,
%``Thermodynamics of Delta resonances,''
Phys.\ Lett.\ B \textbf{433} (1998) 236.
%[nucl-th/9710014].
%%CITATION = NUCL-TH/9710014;%%


\bibitem {denis}
%\cite{Denisenko:1986bb}
%\bibitem{Denisenko:1986bb}
K.~G.~Denisenko and S.~Mrowczynski,
%``Deltas in Hadron Gas,''
Phys.\ Rev.\ C \textbf{35} (1987) 1932.

\bibitem {dick}D.~Parganlija, P.~Kovacs, G.~Wolf, F.~Giacosa and
D.~H.~Rischke,
%``Meson vacuum phenomenology in a three-flavor linear sigma model with (axial-)vector mesons,''
Phys.\ Rev.\ D \textbf{87} (2013) no.1, 014011;
%doi:10.1103/PhysRevD.87.014011
%[arXiv:1208.0585 [hep-ph]];
%%CITATION = doi:10.1103/PhysRevD.87.014011;%%
%\cite{Janowski:2014ppa}
%\bibitem{Janowski:2014ppa}
S.~Janowski, F.~Giacosa and D.~H.~Rischke,
%``Is f0(1710) a glueball?,''
Phys.\ Rev.\ D \textbf{90} (2014) no.11, 114005.
%doi:10.1103/PhysRevD.90.114005
%[arXiv:1408.4921 [hep-ph]].
%%CITATION = doi:10.1103/PhysRevD.90.114005;%%


%\cite{Stachel:2013zma}


\bibitem {begun}
%\cite{Begun:2015ifa}
%\bibitem{Begun:2015ifa}
V.~Begun and W.~Florkowski,
%``Bose-Einstein condensation of pions in heavy-ion collisions at the CERN Large Hadron Collider (LHC) energies,''
Phys.\ Rev.\ C \textbf{91} (2015) 054909;
%doi:10.1103/PhysRevC.91.054909
%[arXiv:1503.04040 [nucl-th]].
%%CITATION = doi:10.1103/PhysRevC.91.054909;%%
%\cite{Begun:2016cva}
V.~Begun,
%``Fluctuations as a test of chemical non-equilibrium at the LHC,''
arXiv:1603.02254 [nucl-th];
%\cite{Begun:2015yco}
%\bibitem{Begun:2015yco}
V.~Begun,
%``High temperature Bose-Einstein condensation,''
arXiv:1512.09157 [nucl-th].
%%CITATION = ARXIV:1512.09157;%%


\bibitem {hiller}
%\cite{Broniowski:2003ax}
%\bibitem{Broniowski:2003ax}
W.~Broniowski, W.~Florkowski and B.~Hiller,
%``Thermal analysis of production of resonances in relativistic heavy ion collisions,''
Phys.\ Rev.\ C \textbf{68} (2003) 034911.
%[nucl-th/0306034].
%%CITATION = NUCL-TH/0306034;%%


\bibitem {milena}T.~Wolkanowski, M.~Soltysiak and F.~Giacosa,
%``$K_{0}^{\ast}(800)$ as a companion pole of $K_{0}^{\ast}(1430)$,''
arXiv:1512.01071 [hep-ph].
%%CITATION = ARXIV:1512.01071;%%
%2 citations counted in INSPIRE as of 14 Mar 2


%\cite{Parganlija:2012fy}


\bibitem {redlichk}
%\cite{Friman:2015zua}
%\bibitem{Friman:2015zua}
B.~Friman et al,
%P. M. Lo, M. Marczenko, K. Redlich and C. Sasaki, ``Strangeness fluctuations from $K-\pi$ interactions,''
Phys.\ Rev.\ D \textbf{92} (2015) no.7, 074003.
%doi:10.1103/PhysRevD.92.074003
%[arXiv:1507.04183 [hep-ph]].
%%CITATION = doi:10.1103/PhysRevD.92.074003;%%

\end{thebibliography}

\documentclass{appolb}%
\usepackage{amssymb}
\usepackage[usenames,dvipsnames]{color}
\usepackage[colorlinks=true,linkcolor=darkblue,citecolor=darkblue,urlcolor=darkblue]%
{hyperref}
\usepackage{graphicx}
\usepackage{makecell}
\usepackage{bm}
\usepackage{amsmath}%
\usepackage{amsfonts}
\pdfoutput=1
\definecolor{darkblue}{RGB}{0,0,196}
\begin{document}

\title{Irrelevance of $f_{0}(500)$ in bulk thermal properties 
\thanks{Talk presented at XI Workshop on Particle Correlations and Femtoscopy, 3-7 November 2015, Warsaw, Poland.}
}
\author{Francesco Giacosa$^{1,2}$, Viktor Begun$^{1}$, \\and Wojciech Broniowski$^{1,3}$
\address{$^{1}$Institute of Physics, Jan Kochanowski University, PL-25406,
Kielce, Poland
\\$^{2}$Institute for Theoretical Physics,
Goethe-University, D-60438, Frankfurt am Main, Germany
\\$^{3}$The
H. Niewodnicza\'{n}ski Institute of Nuclear Physics, Polish Academy of
Sciences, PL-31342, Krak\'{o}w, Poland}
}
\maketitle

\begin{abstract}
We discuss why the scalar-isoscalar resonance $f_{0}(500)$ should in practice
not be included in thermal models describing the freeze-out of heavy-ion
collisions. Its contribution into pion multiplicities is in principle relevant
because it is light and it decays only into pions. However, it is cancelled
to a very good numerical precision by the non-resonant scalar-isotensor
repulsion among pions. Our approach is an application of a well-known theorem
relating spectral function to phase shifts. The numerical results are solely
based on pion-pion scattering data and thus model independent.

\end{abstract}


\section{Introduction}

The scalar-isoscalar resonance $f_{0}(500)$ is now firmly established
\cite{pelaezrev}. The Particle Data Group (PDG) reports the position of its
pole in the range $\left(  400-550\right)  -i(200$-$350$)~\cite{pdg}.
Investigations based on dispersive analysis show even smaller errors: $\left(
400\pm6_{-13}^{+31}\right)  -i(278\pm6_{-43}^{+34}$) in Ref.~\cite{caprini}
and $(457_{-13}^{+14})-i(279_{-7}^{+11})$ in Ref.~\cite{kaminski}.

The resonance $f_{0}(500)$ is the lightest scalar state; moreover, it decays
only into pions. Then, one is lead to think that $f_{0}(500)$ is important for
the determination of pion multiplicities in thermal models for relativistic
heavy-ion collisions (see e.g. Refs.~\cite{wfbook,giorgio} and refs. therein).
Indeed, a simple inclusion of $f_{0}(500)$ as a Breit-Wigner resonance would
lead to a sizable increase (about 3-5\% \cite{andronic}) of pions. However, in
these proceedings (based on the findings of Ref. \cite{our}), we show that
this conclusion is not correct. In fact, when the repulsion of pion-pion
interaction in the scalar-isotensor channel is properly taken into account
using the formalism developed in Refs. \cite{dashen,weinhold,denis}, the
effect of $f_{0}(500)$ cancels to a very good numerical accuracy. We show this
cancellation in a model independent way, since the only input is given by
well-known pion-pion scattering data in these two scalar channels.

\ As a net result, one can neglect in thermal models both the scalar-isoscalar
attraction due to $f_{0}(500)$ and the non-resonant scalar-isotensor repulsion.

\section{Cancellation of $f_{0}(500)$}

A successful description of hadron emissions at the freeze-out of relativistic
heavy-ion collisions is achieved with the help of thermal models.\ For
simplicity, we restrict our presentation to a gas which includes stable pions
($I=0,$ $J^{PC}=0^{-+},$ where $I$ stays for isospin, $J$ for the total spin,
and $P$ and $C$ for parity and charge-conjugation), the $\rho$-resonance
($I=1$, $J^{PC}=1^{--}$ ), the resonance $f_{0}(500)$ ($I=0,$ $J^{PC}=0^{++}%
$), and the non-resonant contribution of the repulsion in the $I=2,$
$J^{PC}=0^{++}$ channel. (Other contributions with different $I$ and $J^{PC}$
correspond to heavier mesons and are neglected here.)

The logarithm of the partition function $Z$ is given by the sum of
contributions of all channels:
\[
\ln Z=\ln Z_{\pi}+\ln Z_{(1,1^{--})}+\ln Z_{(0,0^{++})}+\ln Z_{(2,0^{++}%
)}\text{ .}%
\]
All other thermodynamic quantities follow: $P=-T\ln Z/V,$ $\varepsilon
=-\partial_{\beta}\ln Z/V,$ etc. For what concerns stable pions (we do not
include chemical potentials), one has
\[
\ln Z_{\pi}=3V\int_{p}\ln\left[  1-e^{-\frac{\sqrt{\vec{p}^{2}+M_{\pi}^{2}}%
}{T}}\right]  ^{-1},\text{ }\int_{p}\!=\int\!\frac{d^{3}p}{(2\pi)^{3}}\text{
,}%
\]
where $V$ is the volume, $\vec{p}$ the pion momentum, $M_{\pi}$ the pion mass,
and the factor $3$ the isospin degeneracy. Following Refs.
\cite{dashen,weinhold} we can express the contribution in the channel $I=1,$
$J^{PC}=1^{--}$ as:
\begin{equation}
\ln Z_{(1,1^{--})}=3\cdot3\int_{0}^{\Lambda_{1}}\!\!\!\!\!\!\text{ }%
dm\frac{d\delta_{(1,1)}(m)}{\pi dm}\,\int_{p}\ln\left[  1-e^{-\frac
{\sqrt{p^{2}+m^{2}}}{T}}\right]  ^{-1}\text{ ,} \label{rho}%
\end{equation}
where $\delta_{(1,1)}$ is the measured $\pi\pi$ phase-shift as function of
$m=\sqrt{s}.$ We set $\Lambda_{1}=1$ GeV as maximal energy in the integral,
then only the $\rho$-meson is present in this range. The spectral function of
the $\rho$-meson is related to the phase-shift by \cite{dashen,weinhold}:
\begin{equation}
d_{\rho}(m)=\frac{d\delta_{(1,1)}(m)}{\pi dm}\text{ .}%
\end{equation}
Thus, one can take into account the $\rho$-meson in a thermal gas in a model
independent way by introducing the well-known scattering data in Eq.
(\ref{rho}). For a small width $d_{\rho}(m)$ can be well approximated by a
Breit-Wigner function, $d_{\rho}(m)\simeq\frac{\Gamma}{2\pi}\left[
(m-M_{\rho})^{2}+\Gamma^{2}/4\right]  ^{-1}$ , and, in the limit of zero
width, one correctly obtains $d_{\rho}(m)=\delta(m-M_{\rho}).$ Thus, the
example of the $\rho$-meson shows quite general features of a thermal gas. The
approximation of using a Breit-Wigner expression -typically used in practice- emerges.

We now turn to the main topic of the present work. For $I=J=0,$ the
contribution of $f_{0}(500)$ is included in the integral%
\begin{equation}
\ln Z_{(0,0^{++})}=\int_{0}^{\Lambda_{0}}\!\!\!\!\!\!dm\frac{d\delta_{(0,0)}%
}{\pi dm}\int_{p}\ln\left[  1-e^{-\frac{\sqrt{p^{2}+m^{2}}}{T}}\right]  ^{-1}%
\end{equation}
where $\Lambda_{0}\simeq0.8$ GeV (far above the average mass of $f_{0}(500)$
but below $f_{0}(980),$ which is not considered here). The spectral function
of the $f_{0}(500)$ is given by $d_{f_{0}(500)}(m)=\frac{1}{\pi}%
d\delta_{(0,0)}/dm$. The form of $d_{f_{0}(500)}(m)$ is far from being a
Breit-Wigner \cite{our} and is even not normalized to unity. This is in
agreement with the fact that the resonances $f_{0}(500)$ is not the chiral
partner of the pion and is not a quark-antiquark field \cite{pelaezrev} (the
chiral partner of $\pi$, the `true' $\sigma$ of linear Sigma Models, should be
identified with the heavier scalar resonance $f_{0}(1370)$ \cite{dick}).

As a last step, we consider the joint contribution of both $I=0$ and $I=2$
channels:%
\begin{equation}
\ln Z_{(0,0^{++})}+\ln Z_{(2,0^{++})}=\int_{0}^{\Lambda_{0}}\!\!\!\!\!\!\text{
}dm\left[  \frac{d\delta_{(0,0)}}{\pi dm}+5\frac{d\delta_{(2,0)}}{\pi
dm}\right]  \,\int_{p}\ln\left[  1-e^{-\frac{\sqrt{p^{2}+m^{2}}}{T}}\right]
^{-1}%
\end{equation}
where the factor $5$ in front of $d\delta_{(2,0)}/dm$ is the degeneracy
$2I+1.$ \textit{Data} on pion-pion scattering show the following peculiar fact
\cite{our}:
\begin{equation}
\frac{d\delta_{(0,0)}}{\pi dm}+5\frac{d\delta_{(2,0)}}{\pi dm}\simeq0\text{
for }2M_{\pi}\leq m\lesssim0.8\text{ GeV,}%
\end{equation}
which is valid to a very good numerical accuracy. Then $\ln Z_{(0,0^{++})}+\ln
Z_{(2,0^{++})}\simeq0$! The contribution of $f_{0}(500)$ cancels.

\section{Conclusions}

In these proceedings we have shown that the resonance $f_{0}(500)$ can in
practice be neglected in all isospin-averaged quantities of a thermal hadronic
gas, e.g. for pion multiplicities. Then, the proton-to-pion puzzle becomes
even stronger, leaving other explanations open~\cite{begun}. On the other
hand, in correlation studies of pion-pair production the cancellation does not
occur, hence $f_{0}(500)$ may play a relevant role~\cite{hiller}.

A similar cancellation occurs for the not yet confirmed $K_{0}^{\ast}(800)$
($I=1/2,J^{PC}=0^{++},$ e.g. Ref. \cite{milena} and refs. therein), whose
contribution is (only partly) compensated by the $I=3/2,J^{PC}=0^{++}$ channel
\cite{our,redlichk}.

\bigskip

\textbf{Acknowledgments:} We thank R. Kami\'{n}ski, W. Florkowski, J. Pelaez,
G. Pagliara, V. Mantovani Sarti, T. Wolkanowski for useful discussions. This
research was supported by the Polish National Science Center, Grant DEC-2012/06/A/ST2/00390.

\end{document}